OpenStudy (anonymous):

Pls. Help me answer this. y”+4y’+4y=0 ; y(0)=1, y’(0)=1

OpenStudy (loser66):

where are you stuck?

OpenStudy (anonymous):

y”+4y’+4y=0 ; y(0)=1, y’(0)=1 r^2 +4r+4=(r+2)(r+2) r=-2 and -2

OpenStudy (anonymous):

y=c1e^-2t +C2e^-2t y=(0)=1 y'(0)=1 y'=-C1e^-2t-2C2e-2t

OpenStudy (anonymous):

is it correct???

OpenStudy (loser66):

hold on, y seems not good to me y = e^(-2t) (C1+C2t)

OpenStudy (loser66):

double root r =-2 give you 2 solutions: \(e^{-2t} ~~and~~ t*e^{-2t}\)

OpenStudy (loser66):

general solution is \[y = e^{-2t}(C_1 +C_2t)\]

OpenStudy (loser66):

got me?

OpenStudy (anonymous):

sorry I'm confused. can you pls. explain.

OpenStudy (anonymous):

where did i got wrong?

OpenStudy (loser66):

OK, to yours, C1 e^(-2t) and C2 e^(-2t) is not correct. Re read the material. To DOUBLE ROOT. the second solution MUST time with t

OpenStudy (loser66):

your characteristic equation gives you r = -2 twice, right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

therefore, the solution are: x1 = C1e^(-2t) ok?

OpenStudy (anonymous):

ok

OpenStudy (loser66):

X2 = C2 *t *e^(-2t) that's the formula,

OpenStudy (loser66):

because if you don't time t, so, x1 = x2? when by definition, the solutions must be linear INdependent. to have it, you must time t to x1 to get x2

OpenStudy (anonymous):

so instead of y'=-C1e^-2t-2C2e-2t it should be like this -C1e^-2t- C2 *t *e^(-2t)?

OpenStudy (loser66):

\[y = C_1e^{-2t}+C_2te^{-2t}\] take derivative, you get what? do it carefully, please

OpenStudy (loser66):

\[y' = -2C_1e^{-2t} +C_2e^{-2t} -2C_2e^{-2t}\] got this part?

OpenStudy (loser66):

\(C_2 te^{-2t}\) take product rule for t and e^(-2t)

OpenStudy (anonymous):

yes i got that part

OpenStudy (loser66):

so, you have both y and y' and then, replace initial condition on them to get C1, C2

OpenStudy (loser66):

 got it?